OpenStudy (anonymous):

solve logarithms: log -m+2=4

OpenStudy (anonymous):

My friend sent me this question and he understands not on how to solve this. I myself know nothing of these sort of problems. If I can help him, he will get me a slushie/Icee. :) I really want one. So, please can you assist us in solving this?

hartnn (hartnn):

We would like to help you, but your question is unclear. Can you describe us, what the first term is ?? (or maybe post a screenshot of the entire question)

OpenStudy (anonymous):

the question is unclear :/

OpenStudy (anonymous):

Okay... give me a second

OpenStudy (anonymous):

All it says is to solve the equation.

hartnn (hartnn):

yeah, we know that you need to solve that equation, but how can we help you if we don't understand what the first term is ? is it just \(\log m\) ? why is that - sign there ?

hartnn (hartnn):

you can also draw out your question, by using 'Draw' tool.

hartnn (hartnn):

ok, so first subtract 2 from both sides to isolate the log term, what would you get ?

OpenStudy (anonymous):

log-m=2

OpenStudy (anonymous):

yes.

hartnn (hartnn):

yes, now use the property that \(\huge \log (a)=b \implies a= 10^b \) so, what about \(\log (-m)=2\) ?

OpenStudy (anonymous):

Okay, umm... This is for a friend of mine. I, myself, do not understand what this is. can you explain?

hartnn (hartnn):

i just did. just replace a with -m and b with 2

hartnn (hartnn):

and if your friend is learning logarithms, he would know this : \(\huge \log (a)=b \implies a= 10^b\)

OpenStudy (anonymous):

This is what my friend wrote. : log(-m)=10(2) log(-m)=20 This is what I write: log(-m)=-2 ==> 10^-2

OpenStudy (anonymous):

.01?

hartnn (hartnn):

log (-m) isn't -2, right ? its +2

OpenStudy (anonymous):

okay.

hartnn (hartnn):

let me try with a drawing too, when base of the logarithm is not specified, its 10 so,  got that ?